OpenStudy (anonymous):

divide and simplify. (-15)/ (3/5)=? type an integer or fraction

OpenStudy (diyadiya):

=-15*5/3 =-25

OpenStudy (saifoo.khan):

\[-15 \times \frac 53\] = \[-5 \times 3=-25\]